# Enhanced sensing of Stark weak field under the influence of Aubry-André-Harper criticality

### v2

@@ IV.1 Operator Fisher Information
In Fig. 1(a), we plot OFI, $F_{cdw}$, versus $h$. In the finite-size system for the weak value of $h$, the $F_{cdw}$ has a similar trend as QFI. Initially, it remains flat in the extended region, but after a certain value of $h$, $F_{cdw}$ gradually diminishes and it becomes system size invariant. From this figure, $F_{cdw}$ increases with increasing system sizes in the extended region. The scaling of OFI with L is shown in Fig. 1(c). The violet circles represent the value of $F_{cdw}$ in the extended region for various system sizes up to 987 and the brown dotted line is the fitting function $F_{cdw}\sim L^{\beta}$ with $\beta=6.2$. The scaling associated with $F_{cdw}$ beats the Heisenberg limit (HL) and also the value of scaling exponent $\beta$ is 6.2 which is higher than the scaling of QFI corresponding to the single-particle pure stark case. Hence, the experimentally realizable CDW observable offers a genuine quantum advantage in the measurement precision. We also consider another observable $\hat{O}_{H_{2}}=\sum_{i}i\hat{c}^{\dagger}_{i}\hat{c}_{i}$, which is essentially the last part of our Hamiltonian. In order to evaluate OFI for the operator $\hat{O}_{H_{2}}$, we need to evaluate the expectation values of the operator $\langle\hat{O}_{H_{2}}\rangle$ and $\langle\hat{O}_{H_{2}}\rangle^{2}$, which is basically $\langle\hat{O}_{H_{2}}\rangle^{2}=\sum_{i,j}ij\hat{c}_{i}^{\dagger}\hat{c}_{i}\hat{c}_{j}^{\dagger}\hat{c}_{j}$. In Fig. 1(b), we depict OFI, $F_{H_{2}}$, against $h$ for various system sizes $L$. For a given system size $L$, the OFI, corresponding to the $\hat{O}_{H_{2}}$ has a similar nature like $F_{cdw}$, i.e., initially for the small value of $h$ the $F_{H_{2}}$ is flat and after a certain value of $h$, it decreases gradually. There it becomes system size independent. However, the absolute value of OFI for $\hat{O}_{H_{2}}$ is greater than $F_{cdw}$. In the flat region, the $F_{H_{2}}$ increases steadily with system size $L$. The scaling of $F_{H_{2}}$ in the extended region is shown in Fig. 1(c). The brown star illustrates the value of $F_{H2}$ at $h=10^{-9}$ with different system sizes $L$ up to 987, and the yellow dotted line is the fitting function of $F_{H_{2}}\sim L^{\beta}$ with $\beta=6.7$. This turns out to be a remarkable result as the scaling exponent for OFI $\hat{O}_{H_{2}}$ saturates to that of the QFI.
V Half-filled system as probe
@@
We now focus on the half-filled case. We consider odd lattice sizes with OBC and the corresponding number of particles is either one of these two options: $n_{f}=(L\pm 1)/2$. We perform a slater determination method for this half-filled case. In order to construct the ground state configuration, we diagonalize the single-particle Hamiltonian and took the first $n_{f}$ eigenvectors in a form of Slater determination. The ground state for half-filled case is given by,
@@
[…]
@@
We compute the ground state $F_{Q}$ and $F_{H_{2}}$ for system sizes up to 987 ($n_{f}=493$), where $n_{f}$ is number of fermions. For the half-filled probe, we abstain from performing an averaging over $\phi$. We set $\phi=0$. In the following, we describe the obtained results. Fig. 2(a1) illustrates the QFI for varying $h$. In a finite-size system, within the weak field region, where $h$ is small, the QFI remains constant, exhibiting steady behavior. However, beyond a certain value of $h$, the QFI begins to fluctuate within a specific range due to the trade-off between AAH and the Stark strength. Eventually, its value decreases and stabilizes again. For the weak value of $h$ in a finite-size system, the $F_{Q}$ increases with system size in the flat region of $F_{Q}$. So, in this region, there is a scaling of $F_{Q}$ with fitting function $F_{Q}\sim L^{\beta}$, where we find the scaling exponent $\beta=6.6$. This is shown in Fig. 2(a2) with the maroon dotted line and the maroon circle is the numerical value of $F_{Q}$ at $h=10^{-9}$. The half-filled stark probe also achieves the super-Heisenberg limit with $\beta=6.6$.
@@
The operator $\hat{O}_{cdw}$ is not ideal because the ground state lacks CDW ordering in the localized phase. This means that, unlike in the single-particle scenario, the OFI for $\hat{O}_{cdw}$ does not exhibit proper scaling with $L$. However, the observable $\hat{O}_{H_{2}}=\sum_{i}i\hat{c}_{i}^{\dagger}\hat{c}_{i}$ does provide an appropriate scaling. The corresponding OFI for the finite size system is shown in Fig. 3(b1). This figure illustrates the OFI as a function of $h$. Initially, for small values of $h$, the OFI remains relatively constant. However, beyond a certain threshold of $h$, the OFI starts to fluctuate. The fluctuations dissipate gradually and eventually stabilize again, similar to the behavior of the QFI. Additionally, the value of the OFI nearly converges to that of the QFI. In the flat region where OFI is not changing with $h$, the scaling of OFI for $h=10^{-9}$ is shown in Fig. 3(b2). The violet circle is the value of OFI at $h=10^{-9}$, and the violet dotted line is the fitting function of $F_{H_{2}}\sim L^{\beta}$, where $\beta$ is 6.6. The scaling of OFI for observable $\hat{O}_{H_{2}}$ is also saturated with the scaling of QFI. From Fig. 3(a1) and Fig. 3(b1), it is clear that the scaling sustains over an extended region, which is an added advantage over many quantum critical sensors for which quantum advantage is found over a very narrow region near the quantum criticality. For the half-filled system as a probe, the scaling of QFI beats the Heisenberg limit and gives the super Heisenberg scaling of 6.6. Also, for the half-filled probe, we got an observable that gives a super-Heisenberg scaling of $\beta=6.6$ saturating the scaling of QFI.
VI Conclusions

### v3

@@ IV Single-particle ground state as probe
We first focus on the operator $\hat{O}_{cdw}=\sum_{i}(-1)^{i}\hat{c}_{i}^{\dagger}\hat{c}_{i}$, which measures the occupation imbalance between even and odd sites. This operator reveals the charge-density-wave (CDW) order in a quantum state and can be directly measured in optical lattice setups with ultracold atoms [72, 73, 74]. The computation of OFI is followed the Eq. (2), which entails the average of $\langle O_{cdw}\rangle$ and $\langle\hat{O}_{cdw}^{2}\rangle$, where, $\hat{O}_{cdw}^{2}=\sum_{i,j}(-1)^{i+j}\hat{c}_{i}^{\dagger}\hat{c}_{i}\hat{c}_{j}^{\dagger}\hat{c}_{j}$. In Fig. 1(a), we plot OFI, $F_{cdw}$, versus $h$. In the finite-size system for the weak value of $h$, $F_{cdw}$ has a similar trend as QFI (the QFI plot is shown in Fig. 9 of Ref. [85]). Initially, it remains flat in the extended region, but after a certain value of $h$, $F_{cdw}$ gradually diminishes and it becomes system-size invariant. From this figure, $F_{cdw}$ increases with increasing system size in the extended region. The scaling of OFI with $L$ is shown in Fig. 1(c). The violet circles represent the value of $F_{cdw}$ in the extended region for various system sizes up to 1597, and the dotted violet line is the fitting function $F_{cdw}\sim L^{\beta}$ with $\beta=6.2$. The scaling thus beats the HL. Setting the control parameter at $V=2$, the scaling exponent of OFI provides high precision for estimation of the weak field. Hence, the experimentally realizable CDW observable offers a genuine quantum advantage in the measurement precision.
Next, we consider another observable $\hat{O}_{H_{2}}=\sum_{i}i\hat{c}^{\dagger}_{i}\hat{c}_{i}$, which is essentially the last part of our Hamiltonian. In order to evaluate OFI for the operator $\hat{O}_{H_{2}}$, we need to evaluate the expectation values of the operator $\langle\hat{O}_{H_{2}}\rangle$ and $\langle\hat{O}_{H_{2}}\rangle^{2}$, which is basically $\langle\hat{O}_{H_{2}}\rangle^{2}=\sum_{i,j}ij\hat{c}_{i}^{\dagger}\hat{c}_{i}\hat{c}_{j}^{\dagger}\hat{c}_{j}$. In Fig. 1(b), we depict OFI, $F_{H_{2}}$, against $h$ for various system sizes $L$. For a given system size $L$, the OFI, corresponding to $\hat{O}_{H_{2}}$ has a similar nature to $F_{cdw}$, i.e., initially for a small value of $h$ the $F_{H_{2}}$ is flat and after a certain value of $h$, it gradually decreases. There, it becomes system size independent. However, the absolute value of OFI for $\hat{O}_{H_{2}}$ is greater than $F_{cdw}$. In the flat region, the $F_{H_{2}}$ increases steadily with the system size, $L$. The scaling of $F_{H_{2}}$ in the extended region is shown in Fig. 1(c). The brown star illustrates the value of $F_{H2}$ at $h=10^{-9}$ with different system sizes $L$ up to 1597, and the dotted brown line is the fitting function of $F_{H_{2}}\sim L^{\beta}$ with $\beta=6.7$. This turns out to be a remarkable result as the scaling exponent for OFI $\hat{O}_{H_{2}}$ saturates to that of the QFI.
Resource analysis by taking account of state preparation time. In criticality-based quantum sensing using adiabatic state preparation, it is essential to take into account the preparation time $t$ required to reach the target eigenstate, leading to a modified figure of merit defined by the normalized QFI, $F_{Q}/t$. Near a quantum critical point, the adiabatic evolution time scales with the system size as $t\sim L^{z}$. The exponent $z$ is related to the closing of the energy gap near criticality via $\Delta E\sim L^{-z}$, where $\Delta E$ denotes the energy difference between the ground and first excited states [7]. At the AAH critical point ($V=2$), the energy gap exhibits the scaling $\Delta E\sim L^{-2.37}$ [85], implying $z\approx 2.37$. Incorporating this scaling into the resource cost of adiabatic preparation modifies the QFI scaling from $F_{Q}\sim L^{\beta}$ to $F_{Q}/t\sim L^{\beta-z}$. For our case, we find $F_{Q}/t\sim L^{4.3}$. A similar analysis for the OFI yields $F_{\text{cdw}}/t\sim L^{3.8}$ and $F_{H_{2}}/t\sim L^{4.3}$. These results demonstrate that, even after accounting for the additional cost of state preparation time, both QFI and OFI retain super-Heisenberg scaling, thereby maintaining a strong quantum advantage over classical sensing limits.
Figure 2: Single-party mideigenstate. (a) QFI for the mideigenstate against $h$ for system size $L=144$ (blue), 233 (red), 377 (turquoise), and 610 (black). (b) presents OFI, $F_{cdw}$, for observable $\hat{O}_{cdw}$ with respect to $h$ for system size $L=144$ (blue), 233 (red ), 377 (turquoise), and 610 (black). (c) OFI, $F_{H_{2}}$ for observable $\hat{O}_{H_{2}}$ with respect to $h$ for system size $L=144$ (blue), 233 (red ), 377 (turquoise), and 610 (black). (d) illustrates the scaling of QFI (circular) and OFI, $F_{cdw}$ (square) and $F_{H_{2}}$ (triangular up) for $L=21$, 34, 55, 89, 144, 233, 377, 610, 987, and 1597. The straight lines are the best fits. Whereas $F_{cdw}$ scales as $F_{cdw}\sim L^{5}$ shown by the dotted magenta line, $F_{H_{2}}$ saturates the scaling of the QFI, i.e., $F_{H_{2}}\sim L^{5.6}$, shown by the dotted green line. A configuration averaging over $\phi$ is performed with 8000 random samples.
V Midspectrum
In disordered or tilted (Stark) lattices, the transition between localized and extended behavior does not arise from a single ground-state critical point, but rather from the structure of eigenstates across the spectrum. In the presence of a mobility edge, the localization length diverges as near, the criticality and the rate of divergence is determined by the scaling exponent $\nu$, which is actually energy dependent. This is fundamentally different from a conventional second-order quantum phase transition (QPT), where only the ground state and low-lying excitations capture the QPT directly. In localized systems, each energy window can probe a different effective critical theory, leading to explicitly energy-resolved scaling behavior. Hence, it is interesting to check the effectiveness of the mid-spectrum states as the probe states for high-precision sensing in the localization-based quantum critical sensors. Moreover, the single-particle Stark model can be realized in optical lattices by applying a controlled linear potential, e.g, magnetic field gradients. Transition to an excited state may then be coherently addressed using RF fields, or far-detuned two-photon Raman coupling with negligible spontaneous-emission–induced decoherence.
We investigate the QFI and OFI for midspectrum eigenstates, as depicted in Fig.2. Similar to the ground state, the QFI of the midspectrum eigenstate, $F_{Q}$, exhibits a plateau up to a certain value of the parameter $h$. Beyond this value, $F_{Q}$ begins to decrease nonuniformly. Within the plateau region, the QFI increases with the system size, as demonstrated for $L=144$ (blue), $233$ (red), $377$ (Turquoise), and $610$ (green). This plateau region corresponds to the extended phase, where $F_{Q}$ exhibits a scaling relation $F_{Q}\sim L^{5.6}$, as shown in Fig. 2(a). Although the scaling exponent assumes a lower value in comparison to the ground-state case, it still surpasses the Heisenberg limit for parameter estimation of $h$. In addition, we also analyze the OFI for the CDW operator in the mid-spectrum regime, as shown in Fig.  2(b). Similar to the QFI, the OFI for the CDW operator remains flat within the extended phase for smaller values of $h$. However, beyond this region, it begins to decrease nonuniformly due to the interplay between the two competing potentials, the AA potential and the Stark potential. Within the flat region, the OFI for the CDW operator scales as $F_{cdw}\sim L^{5}$, which also exceeds the Heisenberg limit in measurement precision. Furthermore, Fig. 2(c) illustrates the OFI for the operator $H_{2}$, which is found to saturate the QFI for the midspectrum eigenstate. The scaling relationships for these quantities are summarized in Fig. 2(d). Here, the green circles represent the QFI, while the red triangles correspond to the OFI of $H_{2}$, which saturates the QFI. The green dotted line represents the fitting function with the scaling relation $F_{Q}\sim F_{H_{2}}\sim L^{5.6}$. Additionally, the magenta squares correspond to the OFI for the CDW operator, with a dotted magenta line indicating the fitting function $F_{cdw}\sim L^{5}$. These results collectively highlight the superior scaling behavior of both QFI and OFI in the extended phase, surpassing the Heisenberg limit.
VI Thermal probe
In the many-body scenario, accessing all individual eigenstates of the system becomes computationally and experimentally challenging due to the vast size of the Hilbert space and the complexity of interactions. In practice, the thermal state is often utilized as a probe. The thermal state is defined as
$\rho(h,T)=e^{-H/{k_{B}T}}/\text{Tr}[e^{-H/{k_{B}T}}],$ (4)
where $T$ is the temperature of the system and $k_{B}$ is the Boltzmann constant. In this context, temperature $T$ introduces an additional parameter that influences the behavior of the system. To quantify the sensitivity of the thermal state we have used the definition of QFI as [12]
$F_{Q}(T,h)=\sum_{r,s=1}^{L}\frac{2\text{Re}(\langle\lambda_{r}|\partial_{h}\rho|\lambda_{s}\rangle\langle\lambda_{s}|\partial_{h}\rho|\lambda_{r}\rangle)}{\lambda_{r}+\lambda_{s}},$ (5)
where $\rho$ is the density matrix representing the thermal state and can be written in its spectral decomposition as $\rho=\sum_{r=1}^{L}\lambda_{r}|\lambda_{r}\rangle\langle\lambda_{r}|$. Here $\lambda_{r}$ are the eigenvalues, and $|\lambda_{r}\rangle$ are the corresponding eigenvectors of $\rho$. Re(.) represents the real part of the quantity within the parentheses and the sum excludes the term for which $\lambda_{r}+\lambda_{s}=0$.
Figure 3: Single party thermal state probe for QFI. (a) presents an analysis of the QFI for the thermal state, $\rho(T,h)$, where $T$ stands for the temperature. Setting $h$ at a low value ($h=10^{-8}$), the $F_{Q}$ exhibits distinct values for different system sizes at low temperatures. At sufficiently low temperature, the QFI remains flat and recovers the scaling results obtained from the ground state, i.e., $F_{Q}\sim L^{6.7}$ (see the inset). As the temperature increases, there is a transition in the behavior of $F_{Q}$, and it is marked with a monotonic decay with the temperature. There, the scaling relation turns out to be $F_{Q}\sim T^{-2}$. (b) depicts $F_{Q}$ with temperature $T$ when $h$ is set at a comparatively large value, $h=0.05$, where the considered finite-size systems are in the localized phase at the low temperatures. Here at sufficiently low temperatures, all $F_{Q}$ turns out to be system-size invariant in a similar fashion reported for the ground state. Beyond a transition temperature, $F_{Q}$ decreases with $T$, $F_{Q}\sim T^{-2}$. In both cases, the considered system sizes are $L=55$ (solid), $89$ (dot), $144$ (dash), $233$ (dash-dot), $377$ (dash-dash-dot). A configurational averaing with 500 random realizations of $\phi$ is performed.
Figure 4: Single party thermal state probe for OFI. (a) presents behavior of the OFI, $F_{H_{2}}$, corresponding to the operator, $\hat{O}_{H_{2}}$ in the thermal state $\rho(T,h)$, for $h=10^{-8}$. OFI has a trend similar to QFI (See. Fig. 3), and the scaling relation here is $F_{H_{2}}\sim L^{6.7}$ (see inset) at sufficiently low-temperatures, similar to the behavior observed in the ground state. As the temperature increases, $F_{H_{2}}$ transitions to a decay regime, following the scaling relation $F_{H_{2}}\sim T^{-2}$. (b) Temperature dependence of $F_{H_{2}}$ in the localized phase ($h=0.05$). At low temperatures, $F_{H_{2}}$ collapses across system sizes, resembling ground-state behavior. Beyond a certain temperature, $F_{H_{2}}$ decays with $T$, while maintaining the scaling, $F_{H_{2}}\sim T^{-2}$. (c) For $h=10^{-8}$), $F_{\text{cdw}}$ also follows the ground-state scaling relation, $F_{\text{cdw}}\sim L^{6.2}$ (see inset), at sufficiently low temperatures as well. Above a certain temperature, however, $F_{\text{cdw}}$ decreases with increasing $T$. (d) In the localized phase ($h=0.05$), the temperature dependence of $F_{\text{cdw}}$ shows a collapse across system sizes at low temperatures, consistent with ground-state behavior. At higher temperatures, $F_{\text{cdw}}$ decreases with increasing $T$. In all the cases, considered system sizes are $L=55$ (solid), $89$ (dot), $144$ (dash), $233$ (dash-dot) and $377$ (dot-dash-dot). We have performed an average over 500 random values of $\phi$.
We compute the QFI with respect to the parameter $h$, while temperature $T$ is treated as the varying parameter. To explore the system’s behavior under different conditions, we consider two distinct scenarios, one where the system is in the extended phase at sufficiently low temperatures, for which we set the gradient field as characterized by $h=10^{-8}$, and another where it is in the localized phase, the chosen value for which is $h=0.05$. At low temperatures, specifically when the thermal energy scales as $k_{B}T$, and becomes smaller than the energy gap in the eigenenergy spectrum, $\Delta E$, i.e, $k_{B}T<\Delta E$, the system predominantly behaves like its ground state. This occurs because the ground state possesses the maximum Gibbs probability in the thermal distribution. For $h=10^{-8}$, the sufficiently low-temperature regime exhibits a distinct separation of QFI curves for different system sizes $L$ as depicted in Fig. 3(a). In this regime, the QFI inherits the same scaling behavior as the ground state in the extended phase, i.e., $F_{Q}\sim L^{6.7}$. In contrast, when the system resides in the localized phase ($h=0.05$), the low-temperature QFI plots collapse onto a single curve, irrespective of system size, shown in Fig. 3(b). This collapse is characteristic of the localized systems for the ground state.
For higher temperatures, when $k_{B}T\geq\Delta E$, the system displays intriguing behavior in both the extended and localized phases. As shown in Fig. 3, the QFI monotonically decays as a function of temperature. The QFI decays with temperature as $F_{Q}\sim T^{-2}$. Notably, despite the temperature-induced decay, the QFI in this region continues to scale as $F_{Q}\sim L^{2}$, thereby saturating the Heisenberg limit. This scaling saturation implies that the system’s quantum feature remains robust even at elevated temperatures. To encapsulate this behavior, the QFI can be universally described by the scaling relation
$F_{Q}(T,h)\sim f(h)T^{-2}L^{2},$ (6)
where $f(h)$ is an arbitrary function of $h$. This universal scaling relation offers a comprehensive framework for understanding the interplay between $T$, $h$, and $L$ in determining the QFI.
We also analyze the OFI corresponding to the operator $\hat{O}_{H_{2}}$ for thermal states and observe that its behavior closely resembles that of the QFI with respect to temperature at sufficiently low temperatures. In Fig. 4(a), we present the variation of $F_{H_{2}}$ as a function of temperature $T$ for $h=10^{-8}$, which corresponds to the extended phase of the system. At very low temperatures, $F_{H_{2}}$ plots exhibit distinct separations for different system sizes $L$. This separation is consistent with the behavior previously observed for the ground state, and the scaling in this regime is $F_{H_{2}}\sim L^{6.7}$, indicating a consistent behavior when compared to the ground state case. As the temperature increases, the OFI undergoes a power-law decay following the relation $F_{H_{2}}\sim T^{-2}$. In this relatively high-temperature regime, however, the system exhibits a different scaling behavior– the OFI grows with system size as $F_{H_{2}}\sim L^{2}$, indicating the HL. This highlights the robustness of Heisenberg scaling for thermal states at high temperatures in the extended phase.
In Fig. 4(b), we examine the behavior of $F_{H_{2}}$ in the localized phase, setting $h=0.05$. At low temperatures, $F_{H_{2}}$ for different system sizes collapses onto a single universal curve. This collapse reflects the results in the localized phase corresponding to the ground state case. As the temperature increases, $F_{H_{2}}$ again decays monotonically, following the same power-law relation $F_{H_{2}}\sim T^{-2}$ as in the extended phase. In the high-temperature regime, despite being in the localized phase, the OFI still shows scaling with system size as $F_{H_{2}}\sim L^{2}$, once again reaching the HL.
The behavior of the OFI-corresponding CDW operator is presented in Fig. 4(c). In the delocalized phase, $F_{\text{cdw}}$ increases with system size $L$, exhibiting a scaling exponent of $6.2$, which coincides with that of the ground state. This indicates that at low temperatures, $F_{\text{cdw}}$ retains the same scaling behavior as in the ground state. As the temperature increases, however, $F_{\text{cdw}}$ begins to decrease with $T$, and in this high-temperature regime, the scaling with $L$ is no longer observed. In contrast, in the localized phase [Fig. 4(d)], $F_{\text{cdw}}$ shows a collapse across system sizes at low temperatures, again consistent with ground-state behavior. Upon further increasing the temperature, $F_{\text{cdw}}$ decreases monotonically with $T$, exhibiting HL scaling with system size.
Thus, for relatively large temperatures, the behavior of the OFI across both phases can be captured by a unified scaling ansatz: $F_{H_{2}}(T,h)\sim f(h)T^{-2}L^{2}$, where $f(h)$ is an arbitrary function of $h$. This scaling behavior emphasizes that the value of $h$ strongly affects the low-temperature response of the system. In the relatively high-temperature regime, despite the fact that the system is dominated by thermal fluctuations, we find the Heisenberg-limited scaling with respect to the system size. However, it is worth emphasizing, as that is expected, the absolute values of QFI and OFIs in the high-temperature regime are lower than in the low-temperature regime, implying higher effectiveness of the probes at lower temperatures. For this system, we obtain a quantum advantage in estimating the parameter at low temperatures. As the temperature increases, the advantage in measurement precision gradually decreases. Nevertheless, it is important to note that this trend is not universal. Several alternative models demonstrate that, under specific conditions, high-temperature regimes can actually provide greater precision than their low-temperature counterparts. In such cases, thermal excitations may enhance sensitivity to the parameter of interest [93, 94, 95].
Figure 5: Half-filled case QFI. (a1) QFI for half filled case with respect to $h$. We have taken system size $L(n_{f})$, where $L$ is the system size and $n_{f}$ is the number of filling. This figure shows that in the flat region, the QFI is increasing with system sizes, $L$ = 55 (28) (blue dash-dash), 89 (44) (red dot-dash), 233 (116) (black dot-dot-dash), 377 (188) (green dash-dash-dot). After a certain value of $h$, the QFI decays with certain initial fluctuations, and again it shows steady behavior. (a2) In this plot, we have shown the scaling of QFI with system size $L(n_{f})$. The circle point is numerical value of QFI for $h=10^{-9}$ for $L$ = 21 (10), 55 (28), 89 (44), 233 (116), 377 (188), 987 (493), 1597(798) and the dotted line is the best fitting with fitting function $F_{Q}(h=10^{-9})\sim L^{6.6}$.
Figure 6: Half-filled case OFI. (b1) The figure illustrates the behavior of the OFI for system size $L(n_{f})$, where $n_{f}$ is the number of fermions, with respect to $h$. In the flat region, the OFI increases with system size. However, after reaching a certain value of $h$, the OFI begins to decay with fluctuations and eventually stabilizes into a steady behavior. We have taken $L$ = 55 (blue dash-dash), 89 (red dot-dash), 233 (black dot-dot-dash), 377 ( green dash-dash-dot). (b2) In this plot, we show the scaling of OFI with system size. The violet circular points represent the numerical values of OFI for $h=10^{-9}$ with $L$ = 21 (10), 55 (28), 89 (44), 233 (116), 377 (188), 987 (493), 1597(798), while the violet dotted line represents the best fit. We obtain $F_{H_{2}}(h=10^{-9})\sim L^{6.6}$.
VII Half-filled system as probe
@@
We now focus on the half-filled case. We consider odd lattice sizes with OBC and the corresponding number of particles is either one of these two options: $n_{f}=(L\pm 1)/2$. We perform a slater determination method for this half-filled case. In order to construct the ground-state configuration, we diagonalize the single-particle Hamiltonian and take the first $n_{f}$ eigenvectors in the form of Slater determination. The ground state for the half-filled case is given by,
@@
where $P_{lm}=|\langle\psi_{l}(h)|\psi_{m}(h+dh)\rangle|$. $\psi_{l},\psi_{m}$ are $l\text{'th}$ and $m\text{'th}$ eigenvectors. Now after taylor series expansion, the fidelity can be written as
@@
We compute the ground state $F_{Q}$ and $F_{H_{2}}$ for system sizes up to 987 ($n_{f}=493$), where $n_{f}$ is number of fermions. For the half-filled probe, we abstain from performing an averaging over $\phi$. We set $\phi=0$. In the following, we describe the results that were obtained. Fig. 5(a1) illustrates the QFI for varying $h$. In a finite-size system, within the weak field region, where $h$ is small, the QFI remains constant, exhibiting steady behavior. However, beyond a certain value of $h$, the QFI begins to fluctuate within a specific range due to the trade-off between AAH and the Stark strength. Eventually, its value decreases and stabilizes again. For the weak value of $h$ in a finite-size system, the $F_{Q}$ increases with system size in the flat region of $F_{Q}$. So, in this region, there is a scaling of $F_{Q}$ with fitting function $F_{Q}\sim L^{\beta}$, where we find the scaling exponent $\beta=6.6$. This is shown in Fig. 5(a2) with the dotted maroon line, and the maroon circles denote the numerical value of $F_{Q}$ at $h=10^{-9}$. The half-filled Stark probe also achieves the super-Heisenberg limit with $\beta=6.6$.
@@
The operator $\hat{O}_{cdw}$ is not ideal because the ground state lacks CDW ordering in the localized phase. This means that, unlike in the single-particle scenario, the OFI for $\hat{O}_{cdw}$ does not exhibit proper scaling with $L$. However, the observable $\hat{O}_{H_{2}}=\sum_{i}i\hat{c}_{i}^{\dagger}\hat{c}_{i}$ does provide an appropriate scaling. The corresponding OFI for the finite size system is shown in Fig. 6(b1). This figure illustrates the OFI as a function of $h$. Initially, for small values of $h$, the OFI remains relatively constant. However, beyond a certain threshold of $h$, the OFI starts to fluctuate. The fluctuations dissipate gradually and eventually stabilize again, similar to the behavior of the QFI. Additionally, the value of the OFI nearly converges to that of the QFI. In the flat region where OFI is not changing with $h$, the scaling of OFI for $h=10^{-9}$ is shown in Fig. 6(b2). The violet circle is the value of OFI at $h=10^{-9}$, and the dotted violet line is the fitting function of $F_{H_{2}}\sim L^{\beta}$, where $\beta$ is 6.6. The scaling of OFI for observable $\hat{O}_{H_{2}}$ is also saturated with the scaling of QFI. From Fig. 6(a1) and Fig. 6(b1), it is clear that the scaling sustains over an extended region, which is an added advantage over many quantum critical sensors for which quantum advantage is found over a very narrow region near the quantum criticality. For the half-filled system as a probe, the scaling of QFI beats the Heisenberg limit and gives the super Heisenberg scaling of 6.6. Also, for the half-filled probe, we get an observable that gives a super-Heisenberg scaling of $\beta=6.6$ saturating the scaling of QFI.
VIII Conclusions